OpenStudy (anonymous):

What is the density of hydrogen sulfide (H2S) at 0.6 atm and 337 K? Answer in units of g/L

OpenStudy (hoblos):

use ideal gas law PV=nRT => PV = (m/M)RT => P= dRT/M => d= PM/RT replace the values to get the answer

OpenStudy (anonymous):